OpenStudy (anonymous):

n(2n-1)=k(2k-1)+(4k+1) WHERE , k+1=n, make the right hand side equal to left hand side...... please.....

OpenStudy (mimi_x3):

is that the full question ?

OpenStudy (anonymous):

YES

OpenStudy (mimi_x3):

Like is that the beginning of the question ? since it doesn't make sense ?

OpenStudy (anonymous):

an equation is : n(2n-1)=k(2k-1)+(4k+1) and k+1=n is given, now i want to make right hand side equal to left hand side

OpenStudy (anonymous):

please

OpenStudy (anonymous):

can you understand ?

OpenStudy (mimi_x3):

well, are you sure that its right ? it might be n(2n-1) = n(2k-1) + (4n+1) im not sure.

OpenStudy (mimi_x3):

what does the first question say ?

OpenStudy (mimi_x3):

this is mathematical induction btw

OpenStudy (anonymous):

an equation is given n(2n-1)=k(2k-1)+(4k+1) and it is asking to make the R.H.S. equal to L.H.S. and it is given k+1=n

OpenStudy (anonymous):

now we have to prove R.H.S

OpenStudy (anonymous):

WE DONT HAVE TO TOUCH L.H.S.

OpenStudy (mimi_x3):

well, the question is stated weirdly but anyway (k+1)(2k+1) = k(2k-1)+(4k+1) 2k^2 + 3k + 1 = 2k^2 - k + 4k + 1 2k^2 + 3k + 1 = 2k^2 + 3k + 1 There for RHS=LHS

OpenStudy (anonymous):

no the question is to make: n(2n-1)=n(2n-1)

OpenStudy (mimi_x3):

n(2n-1) = n(2n-1) ??? This is mathematical induction i think, you can just expand and prove it

OpenStudy (anonymous):

this is given : n(2n-1)=k(2k-1)+(4k+1) we have to make it : n(2n-1) = n(2n-1) and a value is given : k+1=n

OpenStudy (anonymous):

?

OpenStudy (mimi_x3):

Hm, well the question is worded weirdly to me, the way i did it was correct to prove RHS=LHS. Hopefully, someoone else would be able to help you sorry.

OpenStudy (anonymous):

no problem , i am very thankfull to you for your kindness and time !!! you tried to help me , this is great to me .....